OpenStudy (anonymous):

I need to prove this equations: 1)sec x - tan x = (cosx / 1+ senx) 2) (1-cos2x / sen 2x) = tanx 3) (cot u -1 / cot u +1) = (1-tan u / 1+ tan u)

OpenStudy (anonymous):

use formula !! try u r self

OpenStudy (anonymous):

thanks for nothing -_-

OpenStudy (anonymous):

\[\sec x - \tan x = \frac{1}{\cos x} - \frac{\sin x}{\cos x}\]\[= \left(\frac{1 - \sin x}{\cos x}\right) \left( \frac{1 + \sin x}{1 + \sin x}\right)\]\[= \frac{1 - \sin ^{2} x}{\cos x \left( 1 + \sin x \right)}\]\[= \frac{\cos^{2} x}{\cos x \left( 1 + \sin x \right)}\]\[= \frac{\cos x}{1 + \sin x}\]Try to do the others by yourself, recall the pythagorean identities and the double-angle formula

OpenStudy (anonymous):

thanks :)